# Comprehensive Analysis of Relative Pressure Estimation Methods Utilizing 4D Flow MRI

### v1

@@ Abstract
4D flow MRI allows for the estimation of three-dimensional relative pressure fields, providing rich pressure information, unlike catheterization and Doppler echocardiography, which provide one-dimensional pressure drops only. The accuracy of one-dimensional pressure drops derived from 4D flow has been explored in previous literature, but additional work must be done to evaluate the accuracy of three-dimensional relative pressure fields. This work presents an analysis of three state-of-the-art relative pressure estimators: virtual Work-Energy Relative Pressure (vWERP), the Pressure Poisson Estimator (PPE), and the Stokes Estimator (STE). Spatiotemporal behavior and sensitivity to noise were determined in silico. Estimators were validated with a type B aortic dissection (TBAD) flow phantom with varying tear geometry and an array of twelve catheter pressure measurements. Finally, the performance of each estimator was evaluated across eight patient cases. In silico pressure field errors were lower in STE compared to PPE, although PPE pressures were less affected by noise. High velocity gradients and low spatial resolution contributed most significantly to local variations in 3D error fields. Low temporal resolution leads to highly transient peak pressure events being averaged, systematically underestimating peak pressures. In the flow phantom analysis, vWERP was the most accurate method, followed by STE and PPE. Each pressure estimator strongly correlated with ground truth pressure values despite the tendency to underestimate peak pressures. Patient case results demonstrated that the pressure estimators could be feasibly integrated into a clinical workflow.
@@
Relative pressure estimation, hemodynamics, 4D flow MRI, aortic dissection
@@ 1 Introduction
Relative pressure differences are important when assessing the severity of a variety of cardiovascular diseases, including aortic stenosis [1], aortic coarctation [2], mitral valve regurgitation [3], and aortic dissection [4]. Until recently, relative pressure differences could only be evaluated by invasive catheterization, which carries inherent risks as a diagnostic tool [5], or by Doppler echocardiography [6]. While Doppler echocardiography is a useful diagnostic tool in a subset of disorders, it measures only one velocity direction. It is, therefore, limited to simple pressure models, such as the simplified Bernoulli equation. 4D flow MRI [7] allows for the measurement of three-dimensional velocity fields over the cardiac cycle, capturing rich blood flow information. This rich flow information enables the estimation of additional hemodynamic variables, such as wall shear stress, kinetic energy, and relative pressure [8, 9, 10]. Pressure estimators that utilize 4D flow MRI have the potential to increase the diagnostic information available by mapping relative pressure across various cardiovascular domains. One-dimensional pressure drops derived from catheter measurements or Doppler echocardiography have been in use for decades, but 4D flow MRI holds the potential for utilizing three-dimensional pressure fields as a diagnostic tool.
@@
We will herein focus on analyzing the performance of pressure estimators in the context of type B aortic dissection (TBAD) [11]. In brief, TBAD refers to a partial thickness entry tear in the descending aorta, creating a second pathway for blood known as the false lumen (FL). The false lumen often reconnects with the true lumen (TL) distally through an exit tear. TBAD contains a wide variety of complex flow patterns and challenging geometries, making it a valuable setting for the evaluation of pressure estimators. TBAD contains regions of low velocities, high velocities, narrow geometries, and highly transient pressure events. Low-velocity regions are helpful when analyzing sensitivity to noise. Narrow domains are essential when analyzing spatial resolution limits; highly transient events are valuable when analyzing temporal resolution limits. Finally, high spatial gradients have been reported to be particularly challenging when estimating hemodynamic parameters [12, 13], meriting further study.
@@
Pressure estimators are highly dependent on the image quality of a given 4D flow acquisition [14]. Consequently, it is crucial to stress-test these estimators under a variety of imaging conditions. Previous work has been done to validate and explore the use of various pressure estimators [15, 16, 17, 18, 14, 19], but additional analysis is necessary. In particular, the spatial distribution of errors requires further study. This work compares CFD-derived pressure fields with estimated pressure fields in a novel voxel-to-voxel analysis. The spatial errors resulting from velocity field corruptions are also determined. Flow phantoms with 12-point arrays of pressure catheter measurements are utilized as ground truth in an in vitro study, enhancing our understanding of the spatial variation of errors in pressure estimations derived from 4D flow acquisitions. In sum, these analyses will add to our understanding of the spatial variations in pressure estimator accuracy rather than the accuracy of one-dimensional pressure drops.
@@
In this paper, the relative pressure estimators Virtual Work-Energy Pressure (vWERP), Pressure Poisson Estimator (PPE), and Stokes Estimator (STE) are analyzed across in silico, in vitro, and in vivo TBAD datasets. For the in silico analysis, we utilize patient-specific CFD simulations to generate synthetic 4D flow MRI as input. The resulting estimated relative pressure fields are compared directly to CFD-derived relative pressure fields. Spatiotemporal and noise sensitivity analyses are performed for both the pressure fields and pressure drops. For the in vitro analysis, we utilize 4D flow acquired across three TBAD flow phantom variations and an array of ground truth pressure catheter measurements [20]. Variations in the flow phantoms allow for a qualitative analysis of the changing pressure fields as the TBAD false lumen tears are modified. Validation of the pressure estimators is performed via comparison to catheter pressure measurements. Finally, pressure drops estimated in eight patient cases are compared across methods for agreement. Sources of error are identified, and each pressure estimator is compared to determine their pros and cons in different use cases.
@@ 2.1 Relative Pressure Estimation Algorithms
Relative pressure estimation algorithms utilize full-field 4D flow MRI velocity measurements as input. These algorithms can be split into fully spatial and plane-to-plane estimators. Fully spatial estimators produce a four-dimensional (3D + time) relative pressure field and include PPE and STE. Plane-to-plane estimators produce a time-resolved relative pressure drop between two user-defined planes (e.g., the pressure drop from the aortic root to the FL exit tear over one cardiac cycle) and include vWERP. The pressure estimation algorithms used herein are derived from the incompressible Navier-Stokes equations [21] given by (1) and (2)
@@ 2.1.1 Pressure Poisson Estimator (PPE)
$\nabla^{2}\hat{p}$ $=-\rho\nabla\cdot({\bf\it u}_{m}\cdot\nabla{\bf\it u}_{m})\quad\text{on }\Omega$ (4)
$\nabla\hat{p}\cdot{\bf\it n}$ $=-\rho({\bf\it u}_{m}\cdot\nabla{\bf\it u}_{m})\cdot{\bf\it n}\quad\text{on }\partial\Omega,$ (5)
@@ 2.1.2 Stokes Estimator (STE)
$\nabla\hat{p}-\mu\nabla^{2}\tilde{{\bf\it u}}$ $=R_{m}\quad\text{on }\Omega$ (6)
@@ 2.2 Implementation of Relative Pressure Estimation Methods
STE, PPE, and vWERP were all implemented using in-house MATLAB codes. All methods utilized a midpoint temporal discretization scheme and second-order spatial derivatives where applicable. The relevant equations for STE and vWERP were discretized with a finite difference method (FDM) that used the measured velocity values directly as input [19]. PPE was discretized using a second-order accurate finite volume method (FVM). For PPE, an FVM was chosen over an FDM to reduce errors introduced by boundary normal vector estimation when prescribing Neumann boundary conditions (5) on the pressure field [26]. Finally, the velocity input was nearest-neighbor upsampled depending on the resolution. The 3.0, 2.0, and 1.5 mm images were upsampled by factors of 4, 4, and 2, respectively. This led to an effective resolution of 0.75, 0.50, and 0.75 mm. All velocity gradients, laplacians, and temporal derivatives were calculated using the raw velocity input before upsampling.
@@ 2.3.1 Model Generation
A patient-specific CFD model of a TBAD patient was utilized as a test bed to analyze PPE, STE, and vWERP. The CFD model allowed for the creation of synthetic 4D flow images at different spatiotemporal resolutions and noise levels to analyze the pressure estimation methods comprehensively. Model geometry was sourced from a clinical CT, segmented in 3D Slicer [27], and meshed with 1.1 mm elements and a 0.9 mm boundary layer in Simmetrix SimModeler [28]. Boundary conditions were prescribed using a 4D flow MRI scan (3T; voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$; temporal resolution = 47 ms, VENC = 2.0 m/s) of the same patient. Inflow and outflow boundary conditions were imposed at the aortic root, brachiocephalic artery, left common carotid artery, left subclavian artery, and true lumen outlet based on flow rates interpolated from 4D flow to 1.0 ms resolution. Flow resistance was estimated by measuring the area of each outlet. The aortic wall was assumed to be completely rigid. A single cardiac cycle was simulated in $\mathcal{C}$Heart [29]. The solved pressure field, $p_{\text{CFD}}$, was projected from the unstructured CFD mesh onto isotropic 0.50 mm and 0.75 mm grids at a 20 ms sampling rate. These projections served as the ground truth.
@@ 2.3.2 Spatiotemporal Analysis
As fully spatial relative pressure estimation methods, PPE and STE could be compared directly with $p_{\text{CFD}}$. Relative error was evaluated at each timestep with (11) and across the entire time domain with (12).
@@
RMSE was also calculated over the entire spatial domain at each timestep and across the entire spatiotemporal domain. In addition to these error metrics, linear regression plots were also generated.
@@ 2.3.3 Noise Sensitivity Analysis
To assess the impact of estimate accuracy, noise fields [31] were added to the projected velocity fields to evaluate the performance of PPE, STE, and vWERP. The noise addition methodology was as follows: Velocity fields were first converted to phase data by specifying a synthetic VENC that was 1.1 times the maximum velocity component across the entire time-resolved velocity field. Synthetic magnitude data was created by multiplying the voxelized CFD mask by a non-zero constant (leaving the non-fluid region equal to 0). The synthetic magnitude and phase images were combined to create complex images that were subsequently transformed to k-space via the fast Fourier Transform. Zero-mean, complex Gaussian noise was added to k-space. The variance of the complex Gaussian was defined to achieve a specified SNR using the equation $\text{SNR}_{\text{db}}=10\log{\frac{P_{x}}{P_{n}}}$ ), where $P_{x}$ is the power of the signal and $P_{n}$ is the power of the noise (which is equivalent to the variance for white Gaussian noise). Finally, the k-space data was transformed back to image space via the inverse Fourier Transform, and the phase images were converted back to velocity images using the previously specified VENC. Twenty-five realizations of SNR = 10 noise fields and SNR = 30 noise fields (Fig. 1c) were calculated for a total of 50 realizations per resolution, resulting in 450 realizations in total. The SNR = $\inf$ was performed implicitly during the spatiotemporal analysis. The same error metrics used in the previous spatiotemporal analysis were used for the noise sensitivity analysis, with some modifications to the correlation analysis — instead of performing a linear regression over one pressure field, linear regressions were performed over all 25 noise realizations at once. Finally, error averages, minimums, and maximums were calculated for each spatiotemporal and noise combination (e.g., the average error was calculated over the 25 realizations of the 1.5 mm $\times$ 40 ms and SNR = 30 combination). Finally, pressure drops were evaluated for their robustness to noise with (13) and (14).
@@ 2.4 Flow Phantoms
PPE, STE, and vWERP were validated with an in vitro flow setup [20]. 4D flow MRI was acquired, and twelve pressure catheter measurements per TBAD tear variation served as ground truth.
@@ 2.4.1 Flow Setup
A full description of the flow phantom can be found in [20] (pump model, exact flow specifications, etc.). A brief overview is provided here. The flow phantom geometry was designed as an embedded compliant 3D-printed TBAD derived from a clinical CT exam. Physiologically accurate pulsatile flow and pressure boundary conditions were applied. 4D flow MRI (3T; voxel size = $1.5\times 1.5\times 1.5$ mm; temporal resolution = 50 ms; VENC = 1.2–1.7 m/s) was acquired across three phantom variations: one baseline phantom (TBADOR) and two additional variations with identical anatomy but with either a reduced size FL entry tear (TBADENT) or a reduced size FL exit tear (TBADEXT) Fig. 2bc. Twelve catheter pressure measurements were acquired at consistent locations across each TBAD tear variation, for a total of 36 unique measurements (Fig. 2a).
@@ 2.4.3 Corruption Analysis
As with the CFD model, the velocity fields were corrupted prior to pressure estimation in order to evaluate each method’s robustness to noise. Noise, however, is already present in the acquired 4D flow, so any additional perturbations are referred to as “corruptions”. Each velocity component ($u_{x},u_{y},u_{z}$) was corrupted with a zero-mean, truncated Gaussian [32] with standard deviation dependent on the level of additional corruption desired ($\Delta\text{SNR}$) according to equation (15).
@@
Each TBAD tear variation was evaluated over 50 realizations with $\Delta$SNR = 30 and $\Delta$SNR = 10, for a total of 300 corrupted realizations per pressure estimation method. Pressure drop deviations were calculated for each unique corrupted velocity field realization. This included measures of mean deviation (13) and max deviation (14). Error averages, minimums, and maximums were recorded across each of the 50 realization combos (e.g., $\text{TBAD}_{\text{ENT}}$ at $\Delta$SNR = 10) and are reported in the supplement.

### v2

# Comprehensive Analysis of Relative Pressure Estimation Methods Utilizing 4D-Flow MRI

@@ Abstract
Magnetic resonance imaging (MRI) can estimate three-dimensional (3D) time-resolved relative pressure fields using 4D-flow MRI, thereby providing rich pressure field information. Clinical alternatives include catheterization and Doppler echocardiography, which only provide one-dimensional pressure drops. The accuracy of one-dimensional pressure drops derived from 4D-flow has been explored previously, but additional work is needed to evaluate the accuracy of 3D relative pressure field estimates. This work presents an analysis of three state-of-the-art relative pressure estimators: virtual Work-Energy Relative Pressure (vWERP), the Pressure Poisson Estimator (PPE), and the Stokes Estimator (STE). The spatiotemporal characteristics and sensitivity to noise were determined in silico. Estimators were then validated using a type B aortic dissection (TBAD) flow phantom with varying tear geometry and twelve catheter pressure measurements. Finally, the performance of each estimator was evaluated across eight patient cases. In silico pressure field errors were lower in STE compared to PPE, although PPE pressures were less noise sensitive. High velocity gradients and low spatial resolution contributed most significantly to local variations in 3D pressure field errors. Low temporal resolution lead to systematic underestimation of highly transient peak pressure events. In the flow phantom analysis, vWERP was the most accurate method, followed by STE and PPE. Each pressure estimator was strongly correlated with ground truth pressure values, despite the tendency to underestimate peak pressures. Patient case results demonstrated that each pressure estimator could be feasibly integrated into a clinical workflow.
@@
Relative pressure estimation, pressure gradient, hemodynamics, 4D-flow MRI, aortic dissection
@@ 1 Introduction
Relative pressure differences are important when assessing the severity of a variety of cardiovascular diseases, including aortic stenosis [1], aortic coarctation [2], mitral valve regurgitation [3], and aortic dissection [4]. Until recently, relative pressure differences could only be evaluated by invasive catheterization, which carries inherent risks as a minor surgery [5], or noninvasively by Doppler echocardiography [6]. While Doppler echocardiography is a useful diagnostic tool in a subset of disorders, it can fail in some patients and only measures one velocity direction. Therefore, it is limited to simple pressure models, such as the simplified Bernoulli equation. 4D-flow MRI [7] measures three-dimensional (3D), time-resolved velocity fields throughout the cardiac cycle, capturing rich blood flow information. This flow information enables the estimation of additional hemodynamic variables, such as wall shear stress, kinetic energy, and relative pressure [8, 9, 10]. Pressure estimators that utilize 4D-flow MRI have the potential to increase the diagnostic information available by mapping relative pressure across various cardiovascular domains. One-dimensional pressure drops derived from catheter measurements or Doppler echocardiography have been in use for decades, but 4D-flow MRI holds the potential for utilizing four-dimensional pressure (3D + time) fields as a diagnostic tool.
@@
Herein, we focus on analyzing the performance of pressure estimators in type B aortic dissection (TBAD) [11]. TBAD refers to a partial thickness entry tear in the descending aorta, creating a second pathway for blood within a false lumen (FL). The false lumen often reconnects with the true lumen (TL) distally through an exit tear. TBAD contains a wide variety of complex flow patterns and challenging geometries, making it a valuable setting for the evaluation of pressure estimators. TBAD contains regions with narrow geometries, highly transient pressure events, and low velocities immediately adjacent to high velocities. Narrow domains are essential when analyzing spatial resolution limits; highly transient events are valuable when analyzing temporal resolution limits. Low-velocity regions are helpful when analyzing sensitivity to noise. Finally, high spatial velocity gradients are particularly challenging when estimating hemodynamic metrics [12, 13], meriting further study.
@@
Pressure estimators are highly dependent on the image quality of a given 4D-flow acquisition [14]. Consequently, it is crucial to stress-test these estimators by perturbing and modifying various image properties (e.g., spatial resolution) and determining their impact on pressure estimation accuracy. Previous work has validated and explored the use of various pressure estimators [15, 16, 17, 18, 14, 19], but additional analysis is necessary – the work above has focused on the accuracy of one-dimensional pressure drops, which don’t take full advantage of the rich velocity information provided by 4D-flow. Furthermore, the spatial distribution of errors requires further study. Herein, we compare CFD-derived pressure fields with estimated pressure fields in a novel voxel-to-voxel analysis. The spatial distribution of errors resulting from synthetic noise is also determined. Compliant TBAD flow phantoms with 12-point arrays of pressure catheter measurements are utilized as ground truth in an in vitro study, enhancing our understanding of the spatial variation of errors in pressure estimations derived from 4D-flow acquisitions. In sum, these analyses will add to our understanding of the spatial variations in pressure estimator accuracy rather than the accuracy of one-dimensional pressure drops.
@@
In this paper, we analyzed three relative pressure estimators: Virtual Work-Energy Pressure (vWERP), the Pressure Poisson Estimator (PPE), and the Stokes Estimator (STE), across in silico, in vitro, and in vivo TBAD datasets. For the in silico analysis, we utilize patient-specific CFD simulations to generate synthetic 4D-flow MRI as input. The resulting estimated relative pressure fields are compared directly to CFD-derived relative pressure fields. Spatiotemporal and noise sensitivity analyses are performed for both the pressure fields and pressure drops. For the in vitro analysis, we utilize 4D-flow acquired across three different compliant TBAD flow phantom with varying FL entry and FL exit tears combined with an array of ground truth pressure catheter measurements [20]. Variations in the flow phantoms allow for a qualitative analysis of the changing pressure fields as the TBAD false lumen tears are modified. Validation of the pressure estimators is performed via comparison to catheter pressure measurements. Finally, pressure drops estimated in eight patient cases are compared across methods for agreement. Sources of error are identified, and each pressure estimator is compared to determine their pros and cons in different use cases.
@@ 2.1 Relative Pressure Estimation Algorithms
Relative pressure estimation algorithms utilize full-field 4D-flow MRI velocity measurements as input. These algorithms can be split into fully spatial and plane-to-plane estimators. Fully spatial estimators produce a four-dimensional (3D + time) relative pressure field and include PPE and STE. Plane-to-plane estimators produce a time-resolved relative pressure drop between two user-defined planes (e.g., the pressure drop from the aortic root to the FL exit tear throughout one cardiac cycle) and include vWERP. The pressure estimation algorithms used herein are derived from the incompressible Navier-Stokes equations [21] given by (1) and (2)
@@ 2.1.1 Pressure Poisson Estimator (PPE)
$\nabla^{2}\hat{p}$ $=\nabla\cdot R_{m}\quad\text{on }\Omega$ (4)
$\nabla\hat{p}\cdot{\bf\it n}$ $=R_{m}\cdot{\bf\it n}\quad\text{on }\partial\Omega,$ (5)
@@ 2.1.2 Stokes Estimator (STE)
$\nabla\hat{p}-\nabla^{2}\tilde{{\bf\it u}}$ $=R_{m}\quad\text{on }\Omega$ (6)
@@ 2.2 Implementation of Relative Pressure Estimation Methods
STE, PPE, and vWERP were all implemented using in-house MATLAB code. All methods utilized a midpoint temporal discretization scheme and second-order spatial derivatives where applicable. The relevant equations for STE and vWERP were discretized with a finite difference method (FDM) that used the measured velocity values directly as input [19]. PPE was discretized using a second-order accurate finite volume method (FVM). For PPE, an FVM was chosen over an FDM to reduce errors introduced by boundary normal vector estimation when prescribing Neumann boundary conditions (5) on the pressure field [26]. Finally, the velocity input was nearest-neighbor upsampled depending on the resolution. The 3.0, 2.0, and 1.5 mm images were upsampled by factors of 4, 4, and 2, respectively. This led to an effective resolution of 0.75, 0.50, and 0.75 mm. All velocity gradients, Laplacians, and temporal derivatives were calculated using the raw velocity input before upsampling.
@@ 2.3.1 Model Generation
A patient-specific CFD model of a TBAD patient was used as a test bed to analyze PPE, STE, and vWERP. The CFD model allowed for the creation of synthetic 4D-flow images at different spatiotemporal resolutions and noise levels to analyze the pressure estimation methods comprehensively. Model geometry was sourced from a clinical CT, segmented in 3D Slicer [27], and meshed with 1.1 mm elements and a 0.9 mm boundary layer in Simmetrix SimModeler [28]. Boundary conditions were prescribed using a 4D-flow MRI scan (3T; voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$; temporal resolution = 47 ms, VENC = 2.0 m/s) of the same patient. Inflow and outflow boundary conditions were imposed at the aortic root, brachiocephalic artery, left common carotid artery, left subclavian artery, and true lumen outlet based on flow rates interpolated from 4D-flow to 1.0 ms resolution. Flow resistance was estimated by measuring the area of each outlet. The aortic wall was assumed to be completely rigid. A single cardiac cycle was simulated in $\mathcal{C}$Heart [29]. The solved pressure field, $p_{\text{CFD}}$, was projected from the unstructured CFD mesh onto isotropic 0.50 mm and 0.75 mm grids at a 20 ms sampling rate. These projections served as the ground truth.
@@ 2.3.2 Spatiotemporal Analysis
PPE and STE produce spatially comprehensive relative pressure estimations ($p_{\text{STE}}$, $p_{\text{PPE}}$) that were compared directly with $p_{\text{CFD}}$. Relative error was evaluated at each timestep with (11) and across the entire time domain with (12).
@@
In addition to the relative error metrics, linear regression plots were generated.
@@ 2.3.3 Noise Sensitivity Analysis
To assess the impact of noise on relative pressure estimate accuracy, random noise fields [31] were added to projected CFD velocity fields before evaluating the performance of PPE, STE, and vWERP. Noise was added as follows: Velocity fields were first converted to phase data by specifying a synthetic VENC that was 1.1 times the maximum velocity component across the entire time-resolved velocity field. Synthetic magnitude data was created by multiplying the voxelized CFD mask by a non-zero constant (leaving the non-fluid region equal to 0). The synthetic magnitude and phase images were combined to create complex images that were subsequently transformed to k-space via the Fast Fourier Transform. Zero-mean, complex Gaussian noise was added to k-space. The variance of the complex Gaussian was defined to achieve a specified SNR using the equation: $\text{SNR}_{\text{db}}=10\log{\frac{P_{x}}{P_{n}}}$ ), where $P_{x}$ is the power of the signal and $P_{n}$ is the power of the noise (which is equivalent to the variance for white Gaussian noise). Finally, the k-space data was transformed back to image space via the inverse Fourier Transform, and the phase images were converted back to velocity images using the specified VENC. Twenty-five realizations of SNR = 10 noise fields and SNR = 30 noise fields (Fig. 1c) were calculated for a total of 50 realizations per resolution, resulting in 450 realizations in total. The SNR = $\inf$ was performed implicitly during the spatiotemporal analysis. The same error metrics used in the previous spatiotemporal analysis were used for the noise sensitivity analysis, with some modifications to the correlation analysis — instead of performing a linear regression over one pressure field, linear regressions were performed over all 25 noise realizations at once. Finally, error averages, minimums, and maximums were calculated for each spatiotemporal and noise combination (e.g., the average error was calculated over the 25 realizations of the 1.5 mm $\times$ 40 ms and SNR = 30 combination). Finally, pressure drops were evaluated for their robustness to noise with (13) and (14). Equations (13) and (14) are referred to in the text as “deviations” because they are designed to measure the spread of the noisy pressure estimations from the baseline (no noise) estimations.
@@ 2.4 Flow Phantoms
PPE, STE, and vWERP were validated with an in vitro flow setup [20]. 4D-flow MRI was acquired, and twelve pressure catheter measurements per TBAD tear variation served as ground truth.
@@ 2.4.1 Flow Setup
A full description of the flow phantom can be found in [20] (pump model, exact flow specifications, etc.). In brief, the flow phantom geometry was designed as an embedded compliant 3D-printed TBAD model derived from a clinical CT exam. Physiologically accurate pulsatile flow and pressure boundary conditions were applied. 4D-flow MRI (3T; voxel size = $1.5\times 1.5\times 1.5$ mm; temporal resolution = 50 ms; VENC = 1.2–1.7 m/s) was acquired across three phantom variations: one baseline phantom (TBADOR) and two additional variations with identical anatomy but with either a reduced size FL entry tear (TBADENT) or a reduced size FL exit tear (TBADEXT) Fig. 2bc. Twelve catheter pressure measurements were acquired at consistent locations across each TBAD tear variation, for a total of 36 unique measurements (Fig. 2a).
@@ 2.4.3 Corruption Analysis
As with the CFD model, the velocity fields were corrupted prior to pressure estimation in order to evaluate each method’s robustness to noise. Noise, however, is already present in the acquired 4D-flow, so any additional perturbations are referred to as “corruptions”. Each velocity component ($u_{x},u_{y},u_{z}$) was corrupted with a zero-mean, truncated Gaussian [32] with standard deviation dependent on the level of additional corruption desired ($\Delta\text{SNR}$) according to equation (15).
@@
Each TBAD tear variation was evaluated over 50 realizations with $\Delta$SNR = 30 and $\Delta$SNR = 10, for a total of 300 corrupted realizations per pressure estimation method. Pressure drop deviations were calculated for each unique corrupted velocity field realization. This included measures of mean deviation (13) and max deviation (14). Error averages, minimums, and maximums were recorded across each of the 50 realization combinations (e.g., $\text{TBAD}_{\text{ENT}}$ at $\Delta$SNR = 10).